A solution of CaCl2 in water forms a mixture that is 36.5% calcium chloride by mass. If the total mass of the mixture is 207.4 g, what masses of CaCl2 and water were used?
1
Expert's answer
2019-09-27T04:45:32-0400

m(CaCl2) = 207.4*36.5% = 75.701 g

Then

m(Water) = m(All) - m(CaCl2) = 207.4 - 75.701 = 131.699 g